I started this and then I could not get the answer to come out.
Evaluate the determinant
5 2 3
5 2 4
2 3 2

Answer by longjonsilver(2297)

Det = 5(2*2 - 4*3) - 2(5*2 - 4*2) + 3(5*3 - 2*2)
Det = 5(4 - 12) - 2(10 - 8) + 3(15 - 4)
Det = 5(-8) - 2(2) + 3(11)
Det = -40 - 4 + 33
Det = -11

jon.